OpenStudy (anonymous):

I'm being asked to calculate some values in a table. I'm given output and TC (total cost) and am asked to calculate AR, TR, MR, AC, MC and profit/loss. I'm also being given the information that this is a monopolist and the market demand function is Q = 10-P. I believe AC is calculated by TC/output, but when trying to calculate the revenue ones, I hit a wall because I don't know the selling price and I'm unsure how to get it out of the market demand function (I know that P = 10-Q)... Also, is MC the incremental addition to AC or TC in this case? Same for MR. Thanks.

OpenStudy (anonymous):

AC=TC/output, correct To calculate revenue, you use the market demand function. Try drawing out the function by choosing all prices from 0 to 10. That will give you the quantity demanded at each price. AR=TR/output TR=output*price MR=the change in TR e.g. if P is 5, Q=5. Therefore, TR is 5*5=25. If we increase output to 6, price= 4. Our new TR is 6*4=24 The change in TR is 25-24=1, therefore MR=1 AC=TC/output MC=change in TC and works the same as MR except you would need to be given a cost function.

OpenStudy (anonymous):

I read you are given TC and output and if tells you TC at each level of output, you can calculate the MC.